Write the Russell-Saunders symbol 2S+!L, for the ground state of the following species: (a)N (b) Zn2+ (c) Crt (d) F (e) V
Answer #1

(a) N = 15², 25² 283 1/170 +1 O -1 me L = Eme = +1 +0-1 = 0 -> 5 S = 32 25+1= 2x}+= J= L+S = 3+ MAIN & &st L, = S+ 17 45 = 45